OpenStudy (anonymous):

Between what two consecutive integers is -√85?

OpenStudy (lucaz):

\[\large{ 8^2=64~~and~~9^2=81 \\ -\sqrt{85}~~is~between~-\sqrt{64}~~and~~-\sqrt{81} }\]

OpenStudy (lucaz):

this means it's between -8 and -9